OpenStudy (anonymous):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 ^3/4 over 2 ^1/2

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

\[\huge\frac{ 2^{\frac{3}{4}}}{2^{\frac{1}{2}}}=2^{\frac{3}{4}-\frac{1}{2}}\]\]

OpenStudy (anonymous):

Not 100% sure how to do that @satellite73